An empty swimming pool measuring 15m by 20m by 1.2m in depth is being filled by taps each discharging water at a rate of 50 litres per minute. If it takes 24 hours to fill the pool, how many taps are used?
Answer by mananth(16946)

swimming pool measuring 15m by 20m by 1.2m
volume of swimming pool = 15*20*1.2
=360m^3
1 m^3=1000 liters
360 m^3= 360000 liters
Let number of taps be x
each minute 50x liters is filled
in 1 hour
the pool fills 50x * 60 liters
= 3000x liters
in 24 hours 24*3000x liters is filled
24*3000x= 360000
x= 360000/(24*3000)
x=5
number of taps required =5